RawlsK29:

Look at the cups shown below (images are not drawn to scale): A cone is shown with width 2 inches and height 4 inches, and a cylinder is shown with width 2 inches and height 6 inches. How many more cubic inches of juice will cup B hold than cup A when both are completely full? Round your answer to the nearest tenth. 8.8 cubic inches 10.1 cubic inches 14.7 cubic inches 22.3 cubic inches

RawlsK29:

1 attachment
RawlsK29:

@Allison @Bearclaws72 @Vocaloid

Allison:

Ahh @Vocaloid

RawlsK29:

Yeah, i am not really sure what the answer is

Allison:

I dunno dis D:

RawlsK29:

Its fine, not a big deal.

Vocaloid:

simply find the volume of cylinder B minus volume of cone A volume of cylinder = pi * r^2 * h volume of cone = (1/3) * pi * r^2 * h